# Optimization of Prompt Learning via Multi-Knowledge Representation for Vision-Language Models

### v2

@@ Abstract
Vision-Language Models (VLMs), such as CLIP, play a foundational role in various cross-modal applications. To fully leverage VLMs’ potential in adapting to downstream tasks, context optimization methods like Prompt Tuning are essential. However, one key limitation is the lack of diversity in prompt templates, whether they are hand-crafted or learned through additional modules. This limitation restricts the capabilities of pretrained VLMs and can result in incorrect predictions in downstream tasks. To address this challenge, we propose Context Optimization with Multi-Knowledge Representation (CoKnow), a framework that enhances Prompt Learning for VLMs with rich contextual knowledge. To facilitate CoKnow during inference, we trained lightweight semantic knowledge mappers, which are capable of generating Multi-Knowledge Representation for an input image without requiring additional priors. Experimentally, We conducted extensive experiments on 11 publicly available datasets, demonstrating that CoKnow outperforms a series of previous methods. We will make all resources open-source: https://github.com/EMZucas/CoKnow.
1 Introduction
@@
Figure 1: Comparison before and after introducing Multi-Knowledge Representation in zero-shot scenarios.
To enhance the performance of CLIP [22] in downstream tasks, many methods have been proposed, roughly categorized into two main types: Fine Tuning and Prompt Learning. To avoid the high computational load of full parameter tuning, Fine Tuning methods typically freeze CLIP’s parameters and introduce trainable additional modules to adapt to downstream tasks. Prompt Learning methods, on the other hand, attempt to learn better contextual templates to replace manually crafted prompts such as a photo of a [CLASS]. These methods effectively improve the performance of CLIP, but there is still much room for improvement in many tasks, especially in few-shot tasks in low-resource scenarios.
The improvement space for pretrained VLMs in downstream tasks can come from the potential they have yet to be elicited. One possible explanation is that a simple text prompt may not adequately represent the complexity of an image, which typically contains more information than its captions. In reality, a single object can be described in various ways, depending on the level of detail or perspective. While models like CLIP have been trained on a vast dataset of image-text pairs and have learned rich relationships between visual and textual content, they may struggle with simple text prompts that fail to capture the full semantic meaning of an image. To fully utilize the capabilities of CLIP, we propose to enhance the prompt context by incorporating knowledge from multiple perspectives at multiple abstraction levels, or in short Multi-Knowledge. In downstream tasks of image recognition, we define the global representation of a piece of text in VLMs as Multi-Knowledge Representation. Specifically, a natural language description consists of multiple words, each represented by a token in VLMs. Each word also represents different pieces of knowledge, and their encoded global representations constitute the Multi-Knowledge Representation.
@@
To leverage Multi-Knowledge Representation and the successful techniques of fine-tuning and soft prompt tuning for VLMs, we propose CoKnow, which consists of two learnable key modules, as shown in Figure 2. The first module is an optimizer for Prompt Learning guided by Multi-Knowledge representation, allowing for adaptive learning of prompt templates rich in domain knowledge for various downstream tasks. The second module is a lightweight semantic knowledge mapper, which, once trained, can directly generate the corresponding Multi-Knowledge Representation from images without additional input. This mapper is designed as a general module that can operate in a plug-and-play style with general VLMs, not limited to CLIP. Notably, CoKnow leverages the Large Language Model GPT-4 [2] as the knowledge generator, automatically producing Multi-Knowledge Representation using a set of simple prompt templates. Experimentally, We conducted experiments on 11 publicly available datasets spanning various downstream domains using CoKnow. The results, surpassing several previous few-shot methods, confirm that CoKnow is an effective approach for VLM applications. Our main contributions are as follows:
  • To enhance VLMs performance in downstream tasks like classification, We propose Multi-Knowledge Representation, including image captions at visual level, non-visual descriptions at abstract level, as well as their combinations.
  • We introduce the CoKnow framework, featuring a trainable prompt optimizer and lightweight visual-text semantic mappers, to support Multi-Knowledge Representation in tasks based on vision-language models.
  • Effective and scalable methods are provided by us for generating knowledge across diverse categories. For 11 publicly available datasets, we construct 6603 distinct knowledge descriptions corresponding to 2201 object categories using GPT-4.
2 Related Work
Table 1: Comparison of image recognition methods from different series for vision-language models. Extra Modules: Refers to incorporating other network modules for training. Soft Prompt: Refers to learning adaptable templates using learnable embedding vectors for downstream tasks. Image Rep: Represents the utilization of image representation. Fixed Template Rep: Represents the utilization of fixed templates from the downstream dataset, such as a photo of a [classname]. Multi-Knowledge Rep: Represents the utilization of multi-knowledge representation. ✓indicates that this part of the content is utilized in the method.
@@ 2 Related Work
Nowadays, significant advancements are being made in vision-language models [13, 23, 25, 31, 32], which create rich multimodal representations between natural language and vision, facilitating a wide array of downstream tasks. Among these models, CLIP [22] stands out as a prominent method. By undergoing self-supervised training on 400 million image-text pairs, CLIP learns joint representations for images and text, showcasing outstanding performance across various downstream visual tasks. Although pre-trained VLMs can represent image-text pairs, applying them effectively to downstream tasks still faces severasl challenges. Recently, many studies have shown improved performance in downstream tasks, such as few-shot image recognition [36, 18, 34], by customizing CLIP through Fine Tuning and Prompt Learning.
2.1 Fine Tuning
@@
In downstream tasks with few samples, full-finetuning [14, 19, 33] of CLIP often leads to overfitting issues, resulting in poor generalization performance on test data. Recently, introducing additional modules to CLIP and fine-tuning them has proven to be an effective method for adapting it to downstream image recognition tasks. CLIP adopts linear probing [22, 4], which involves introducing an additional linear layer for fine-tuning on downstream tasks. CALIP [10] invoked a parameter-free attention mechanism to promote interactions between visual and textual modalities, exploring essential cross-modal informational features. Its parametric iteration, CALIP-F [10], particularly excelled in few-shot testing environments, showcasing outstanding efficiency. WiSE-FT [28] proposed a weight-space ensemble approach, merging CLIP’s pre-trained and fine-tuned weights to bolster robustness against atypical distributions. Houlsby and colleagues [12] introduced an adapter strategy that embeds learnable linear layers into transformer layers while immobilizing the network’s backbone, providing a streamlined path for fine-tuning downstream tasks. Gao et al. [8] formulated the CLIP-Adapter, integrating feature adapters for nuanced fine-tuning within the visual or language branches. Zhang et al. Reference [34] launched the Tip-Adapter, which employs a key-value cache model conceived from a limited-sample training set, enhancing initialization for quicker fine-tuning convergence. Lin et al. [16] ventured into cross-modal few-shot learning by repurposing textual features as training samples during the fine-tuning phase, paving a novel pathway for utilizing textual data in training. These methods adapt CLIP for downstream tasks by introducing additional modules. Similarly, we trained the semantic knowledge mappers, which can automatically generate corresponding Multi-Knowledge Representation without the need for additional input, offering the advantage of plug-and-play functionality.
Figure 2: The overall architecture and training process of CoKnow. The figure illustrates the input of three types of templates into the text encoder: Learnable Context, Multi-Knowledge, and Hand-Crafted Template, when the category is airplane. The output of the image encoder undergoes contrastive loss calculation with the corresponding target template embeddings after passing through the semantic knowledge mappers. The original image representation and the mapped image representation are added with $\beta$ weights before undergoing contrastive loss calculation with the final embedding representation of Learnable Contexts. Throughout the entire process, we only train lightweight semantic knowledge mappers to optimize context learning.
2.2 Prompt Learning
@@
Another major challenge of applying large-scale pre-trained VLMs like CLIP to image recognition tasks is prompt engineering, which demands domain-specific expertise downstream and is extremely time-consuming, requiring substantial time investment for adjustments. To address this issue, CoOp [36] first introduced prompt learning into CLIP, achieving outstanding results in downstream few-shot image recognition tasks. Subsequent improvements have enabled a variety of tasks, including fine-grained object retrieval [27] and image-text classification [9]. Furthermore, CoCoOp [35] enhanced generalization by incorporating image features into each prompt embedding using lightweight networks. ProGrad [37] refined downstream task adaptation by correcting feedback gradients. Recently, KgCoOp [30] proposed constraining learnable prompt embeddings with common knowledge to enhance transfer generalization. These methods optimize learnable templates by leveraging image representations, and further, KgCoOp introduces hand-crafted template representations for optimizing learnable templates.
2.3 Multi-Knowledge Representation
@@
Both existing methods, Prompt Learning and Model Fine-tuning, overlook the potential for incorporating a broader range of domain-specific knowledge representations corresponding to image representations, which can be articulated using natural language. We refer to this global semantic representation described in natural language as Multi-Knowledge Representation. Recently, CLIP-A-self [18] leverages GPT-4 to generate multiple visual description texts as input to the text encoder, and utilizes self-attention mechanism networks to aggregate them on each sentence, achieving good generalization effects. These visual descriptions are remarkably similar to the VK type within Multi-Knowledge. We first propose the use of Multi-Knowledge Representation to optimize learnable contextual templates, delivering domain knowledge information to downstream tasks through adaptive means, while reducing human involvement, thus making the process more convenient. And explored the effectiveness of other types of Multi-Knowledge, including VK, NVK, and PK.
Figure 3: t-SNE visualization. Different colors represent different categories in the CIFAR-10 dataset. In the case of zero-shot, compared to the original CLIP, a clear division into 10 clusters can be observed.
@@
In Table 1, we clearly delineate the distinctions between our method and various others. Existing approaches primarily focus on learning soft prompts and fine-tuning Extra Modules, lacking guidance from Multi-Knowledge Representation. Methods centered around domain knowledge require manual adjustments of this natural language aspect, lacking an adaptive template design approach. Our proposed method, CoKnow, addresses these shortcomings. For context template optimization, we not only leverage representations of images and fixed text but also introduce additional Multi-Knowledge Representation for context optimization, achieving outstanding results in experiments for the first time. Moreover, due to the presence of lightweight semantic knowledge mappers in CoKnow, no additional input is required during the inference stage. Prediction results can be obtained solely by inputting the image and the trained context template.
3 Method
Our method is based on a pre-trained vision-language model CLIP [22], as shown in Figure 2. In terms of model architecture, we only introduce the lightweight semantic knowledge mappers, which can generate Multi-Knowledge Representation without additional priors. The Multi-Knowledge Representation is important. We introduce three different types of Multi-Knowledge Representation, namely VK, NVK, and PK, respectively, on CIFAR-10 [1] dataset. By incorporating the Multi-Knowledge Representation with the corresponding original image representation, we obtain new representation vectors. (Same as the approach in Figure 1(b).) The results are visualized using t-SNE as depicted in Figure 3. (For specific details and additional results, please refer to Appendix.) It can be observed that introducing different types of Multi-Knowledge Representation allows for excellent partitioning of the new representations into 10 clusters. We have constructed a Prompt learning framework guided by Multi-Knowledge Representation. Next, we will introduce the technical details and principles of our approach.
3.1 Text Encoder Input
@@
Hand-Crafted Template  Since the introduction of CLIP, the method of inserting category names into Hand-Crafted Templates for image recognition has been widely adopted by subsequent researchers. In various downstream tasks, the approach of Hand-Crafted Templates has demonstrated superior performance. In CoKnow, we continue to leverage the Hand-Crafted Templates proposed by previous researchers, as shown in Figure 2. This type of hand-crafted template, such as a photo of a [CLASS]. Here, [CLASS] represents the insertable category name.

### v3

@@ Abstract
Vision-language models (VLMs), such as CLIP, play a foundational role in various cross-modal applications. To fully leverage the potential of VLMs in adapting to downstream tasks, context optimization methods such as prompt tuning are essential. However, one key limitation is the lack of diversity in prompt templates, whether they are hand-crafted or learned through additional modules. This limitation restricts the capabilities of pretrained VLMs and can result in incorrect predictions in downstream tasks. To address this challenge, we propose context optimization with multi-knowledge representation (CoKnow), a framework that enhances prompt learning for VLMs with rich contextual knowledge. To facilitate CoKnow during inference, we train lightweight semantic knowledge mappers, which are capable of generating multi-knowledge representations for an input image without requiring additional priors. Experimentally, we conduct extensive experiments on 11 publicly available datasets, demonstrating that CoKnow outperforms a series of previous methods.
Index Terms:
Vision-language models, few-shot learning, prompt learning, multi-knowledge
I Introduction
@@
Figure 1: The result predicted by CLIP without introducing multi-knowledge representation is shown in Figure (a). The result predicted by CLIP after introducing multi-knowledge representation is shown in Figure (b). Introducing multi-knowledge representation improves the prediction accuracy.
To enhance the performance of CLIP [1] in downstream tasks, many methods have been proposed, which can be roughly categorized into two main types: fine tuning and prompt learning. To avoid the high computational load of full parameter tuning, fine-tuning methods typically freeze the CLIP parameters and introduce trainable additional modules to adapt to downstream tasks. Prompt learning methods, on the other hand, attempt to learn better contextual templates to replace manually crafted prompts such as a photo of a [CLASS]. These methods effectively improve the performance of CLIP, but there is still much room for improvement in many tasks, especially in few-shot tasks in low-resource scenarios.
The improvement space for pretrained VLMs in downstream tasks can come from the potential they have yet to be elicited. One possible explanation is that a simple text prompt may not adequately represent the complexity of an image, which typically contains more information than its captions do. In reality, a single object can be described in various ways, depending on the level of detail or perspective. While models such as CLIP have been trained on a vast dataset of image-text pairs and have learned rich relationships between visual and textual content, they may struggle with simple text prompts that fail to capture the full semantic meaning of an image. To fully utilize the capabilities of CLIP, we propose enhancing the prompt context by incorporating knowledge from multiple perspectives at multiple abstraction levels, or in short, Multi-Knowledge. In downstream tasks of image recognition, we define the global representation of a piece of text in VLMs as a multi-knowledge representation. Specifically, a natural language description consists of multiple words, each represented by a token in VLMs. Each word also represents different pieces of knowledge, and their encoded global representations constitute the multi-knowledge representation.
@@
Figure 2: Instances of Multi-Knowledge. These include categories such as airplanes, human faces, ship wheels, pandas, gyoza food, and bubble textures. Here, blue represents the content generated by GPT-4, orange represents the content generated by LLama3.1, and green represents the content generated by Qwen2.
To leverage multi-knowledge representation and the successful techniques of fine-tuning and soft prompt tuning for VLMs, we propose CoKnow, which consists of two learnable key modules, as shown in Figure 4. The first module is an optimizer for prompt learning guided by multi-knowledge representation, allowing for adaptive learning of prompt templates rich in domain knowledge for various downstream tasks. The second module is a lightweight semantic knowledge mapper, which, once trained, can directly generate the corresponding multi-knowledge representation from images without additional input. This mapper is designed as a general module that can operate in a plug-and-play style with general VLMs and is not limited to CLIP. Notably, CoKnow leverages the large language model GPT-4 [8] as the knowledge generator, automatically producing multi-knowledge representations via a set of simple prompt templates. Current large language models have demonstrated remarkable emergent intelligence, generalization capabilities, and knowledge processing abilities [9]. Experimentally, we conducted experiments on 11 publicly available datasets spanning various downstream domains via CoKnow. The results, which surpass several previous few-shot methods, confirm that CoKnow is an effective approach for VLM applications. Our main contributions are as follows:
  • To increase the performance of VLMs in downstream tasks such as classification, we propose multi-knowledge representations, including image captions at the visual level, nonvisual descriptions at the abstract level, and their combinations. Simultaneously, we validate the effectiveness of the multi-knowledge representation.
  • We introduce the CoKnow framework, which features a trainable prompt optimizer and lightweight visual-text semantic mappers, to support multi-knowledge representation in tasks based on vision-language models.
  • We provide effective and scalable methods for generating knowledge across diverse categories. For 11 publicly available datasets, we construct 6603 distinct knowledge descriptions corresponding to 2201 object categories via GPT-4. All resources will be made open source to support future research efforts.
TABLE I: Comparison of image recognition methods from different series for vision-language models. ”✓” indicates that this part of the content is utilized in the method.
@@ II Related work
Currently, significant advancements are being made in vision-language models [16, 17, 18, 19, 20], which create rich multimodal representations between natural language and vision, facilitating a wide array of downstream tasks. Among these models, CLIP [1] stands out as a prominent method. By undergoing self-supervised training on 400 million image-text pairs, CLIP learns joint representations for images and text, showing outstanding performance across various downstream visual tasks. Although pretrained VLMs can represent image-text pairs, applying them effectively to downstream tasks still faces several challenges. Recently, many studies have shown improved performance in downstream tasks, such as few-shot image recognition [13, 12, 11], by customizing CLIP through fine tuning and prompt learning.
II-A Fine Tuning
@@
In downstream tasks with few samples, full-tuning [21, 22, 23] of CLIP often leads to overfitting issues, resulting in poor generalization performance on test data. Recently, introducing additional modules to CLIP and fine-tuning them has proven to be an effective method for adapting it to downstream image recognition tasks. CLIP adopts linear probing [1, 24], which involves introducing an additional linear layer for fine-tuning downstream tasks. CALIP [25] invokes a parameter-free attention mechanism to promote interactions between visual and textual modalities, exploring essential cross-modal informational features. Its parametric iteration, CALIP-F [25] particularly excelled in few-shot testing environments, showcasing outstanding efficiency. WiSE-FT [26] proposes a weight-space ensemble approach that merges CLIP’s pretrained and fine-tuned weights to bolster robustness against atypical distributions. Houlsby and colleagues [27] introduced an adapter strategy that embeds learnable linear layers into transformer layers while immobilizing the network’s backbone, providing a streamlined path for fine-tuning downstream tasks. Gao et al. [10] formulated the CLIP-Adapter, which integrates feature adapters for nuanced fine-tuning within the visual or language branches. Zhang et al. [11] launched Tip-Adapter, which employs a key-value cache model conceived from a limited-sample training set, enhancing initialization for quicker fine-tuning convergence. Lin et al. [28] ventured into cross-modal few-shot learning by repurposing textual features as training samples during the fine-tuning phase, paving a novel pathway for utilizing textual data in training. These methods adapt CLIP for downstream tasks by introducing additional modules. Similarly, we trained the semantic knowledge mappers, which can automatically generate corresponding multi-knowledge representations without the need for additional input, offering the advantage of plug-and-play functionality.
II-B Prompt Learning
@@
Another major challenge of applying large-scale pretrained VLMs such as CLIP to image recognition tasks is prompt engineering, which demands domain-specific expertise downstream and is extremely time-consuming, requiring substantial time investment for adjustments. To address this issue, CoOp [13] first introduced prompt learning into CLIP, achieving outstanding results in downstream few-shot image recognition tasks. Subsequent improvements have enabled a variety of tasks, including fine-grained object retrieval [29] and image-text classification [30]. Furthermore, CoCoOp [14] enhances generalization by incorporating image features into each prompt embedding via lightweight networks. ProGrad [31] refines downstream task adaptation by correcting feedback gradients. Recently, KgCoOp [15] proposed constraining learnable prompt embeddings with common knowledge to enhance transfer generalization. These methods optimize learnable templates by leveraging image representations, and further, KgCoOp introduces handcrafted template representations for optimizing learnable templates.
II-C Multi-Knowledge Representation
@@
Both existing methods, prompt learning and model fine-tuning, overlook the potential for incorporating a broader range of domain-specific knowledge representations corresponding to image representations, which can be articulated via natural language. We refer to this global semantic representation described in natural language as multi-knowledge representation. Recently, CLIP-A-self [12] leverages GPT-4 to generate multiple visual description texts as inputs to the text encoder and uses self-attention mechanism networks to aggregate them on each sentence, achieving good generalization effects. These visual descriptions are remarkably similar to the VK type within Multi-Knowledge. We first propose the use of multi-knowledge representation to optimize learnable contextual templates, delivering domain knowledge information to downstream tasks through adaptive means while reducing human involvement, thus making the process more convenient. We explored the effectiveness of other types of Multi-Knowledge, including VK, NVK, and PK. It is worth noting that we choose the powerful GPT-4 large language model version to generate multi-knowledge to assist our classification task, rather than directly using its multimodal capabilities for image classification. This decision is based on the fact that the multimodal version of GPT-4 still has significant limitations in fine-grained image question-answering tasks, with poor classification performance between fine-grained categories. Additionally, as a VQA model, it heavily relies on manually adjusted prompt settings, leading to inefficient classification and inevitable hallucination problems. In contrast, when GPT-4 has a clear understanding of a particular object category, i.e., when it clearly knows the name of the object, it can provide rich knowledge related to that object. Finally, we present the experimental results of using the GPT-4 Vision version and GPT-4o for direct image classification in Figure 3. These results are compared with those obtained by introducing PK on CLIP-RN50 (Ours). The prompt we use is: You are an image classification AI assistant, please perform category recognition on this image. The classification accuracy is below 30% for most categories.
Figure 3: Image classification is performed on the Oxford Pets dataset using multimodal large models and our method.
Figure 4: The overall architecture and training process of CoKnow. The figure illustrates the input of three types of templates into the text encoder: learnable context, Multi-Knowledge, and hand-crafted template, when the category is airplane. The output of the image encoder undergoes contrastive loss calculations with the corresponding target template embeddings after passing through the semantic knowledge mappers. The original image representation and the mapped image representation are added with $\beta$ weights before performing contrastive loss calculations with the final embedding representation of the learnable contexts. Throughout the entire process, we train only lightweight semantic knowledge mappers to optimize context learning.
Figure 5: The prediction results with the introduction of different Multi-Knowledge. (a) The figure presents the prediction results obtained via GPT-4, Qwen2, Llama3.1 and GPT4-o with a ResNet-50 backbone. (b) The figure illustrates the prediction results achieved with GPT-4, Qwen2, Llama3.1 and GPT4-o utilizing a ViT-B/32 backbone. (c) The figure displays the prediction results derived from GPT-4, Qwen2, Llama3.1 and GPT4-o employing a ViT-B/16 backbone. Among all the results, the Multi-Knowledge approach utilizing GPT-4 results in greater stability.
@@
In Table I, we clearly delineate the distinctions between our method and various others. (Extra Modules refers to incorporating other network modules for training. Soft Prompt refers to learning adaptable templates using learnable embedding vectors for downstream tasks. Image Rep represents the utilization of image representation. Fixed Template Rep represents the utilization of fixed templates from the downstream dataset, such as a photo of a [classname]. Multi-Knowledge Rep represents the utilization of multi-knowledge representation.) Existing approaches focus primarily on learning soft prompts and fine-tuning extra modules, lacking guidance from multi-knowledge representations. Methods centered on domain knowledge require manual adjustments of this natural language aspect, lacking an adaptive template design approach. Our proposed method, CoKnow, addresses these shortcomings. For context template optimization, we not only leverage representations of images and fixed text but also introduce additional multi-knowledge representations for context optimization, achieving outstanding results in experiments for the first time. Moreover, owing to the presence of lightweight semantic knowledge mappers in CoKnow, no additional input is required during the inference stage. The prediction results can be obtained solely by inputting the image and the trained context template.
III Method
Our method is based on a pretrained vision-language model CLIP [1], as shown in Figure 4. In terms of model architecture, we introduce only lightweight semantic knowledge mappers, which can generate multi-knowledge representations without additional priors. The multi-knowledge representation is crucial. To demonstrate its effectiveness, we introduce three different types of multi-knowledge representations-VK, NVK, and PK-on the CIFAR10, CIFAR100, and STL10 datasets, respectively. By combining the multi-knowledge representation with the corresponding original image representation, we obtain new representation vectors (the same approach as in Figure 1(b)). In Figure 5, we present many experimental results, including various backbone visual language models, such as ResNet-50, ViT-B/32, and ViT-B/16, as well as different large language models (LLMs), including GPT-4, Qwen2 [32], and Llama3.1 [33]. These results indicate that the embedded representation of Multi-Knowledge can significantly enhance the overall predictive accuracy. However, when different LLMs are compared, the quality of Multi-Knowledge generation varies across models with different capabilities, leading to inconsistencies in predictive performance. For example, on the CIFAR10 [7], CIFAR100 , and CUB-200-2011 [34] datasets, the multi-knowledge generated by each large language model improves the prediction accuracy of each backbone network. However, on the STL10 dataset, except for the multi-knowledge generated by GPT-4, which consistently enhances prediction accuracy, neither Qwen2 nor Llama3.1 improves accuracy across all types of multi-knowledge. At the same time, the knowledge quality generated by the powerful reasoning model GPT4-o is unstable and does not perform well on every dataset.
Furthermore, as shown in Figure 5, we perform a t-SNE visualization of the combined vector results from the original image representation and the multi-knowledge representation (Multi-Knowledge generated by GPT-4). On various datasets, these representations are clearly segmented into multiple clusters, corresponding to different categories. We construct a prompt learning framework guided by multi-knowledge representation. Next, we introduce the technical details and principles of our approach.
III-A Text Encoder Input
@@ III-A1 Hand-Crafted Template
Since the introduction of CLIP, the method of inserting category names into hand-crafted templates for image recognition has been widely adopted by subsequent researchers. For various downstream tasks, the hand-crafted template approach has demonstrated superior performance. We continue to leverage the hand-crafted templates proposed by previous researchers, as shown in Figure 4. This type of hand-crafted template is a photo of a [CLASS]. Here, [CLASS] represents the insertable category name.
III-A2 Multi-Knowledge